OpenStudy (anonymous):

A rectangle has width 22 in and length 18 in. Find the length of another rectangle of equal area, if the width is 20 in. A. 16.4 in B. 17.6 in C. 19.8 in

OpenStudy (anonymous):

396= 20x x= 19.8

OpenStudy (anonymous):

22 * 18 = 20 * l Solve for l

OpenStudy (anonymous):

C. \[\frac{99}{5}=19.8 \]

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

Area of Rectangle= length * width =22*18=396 Area= 20 * length length= area/20 =19.8

OpenStudy (anonymous):

C